Grade 11General Physics

the frequency of first overtone of a closed organ pipe of length L is f1. a hole is made at a distance L/6 from the closed end. now the frequency of first overtone of open pipe is f2. then ratio of f1/f2 is ?
….…..the problem is i don’t understand how will the wavelength of sound will change when hole is made ….….well the is 5/8 ….….

shashank
11 Years agoGrade 11

1 Answer

Abhishek Kumar
11 Years ago
When the hole is made, the hole acts as the pressure node or the displacement antinode as the open end of the tube acts.

f1 = 3V/4L
When, opening is made at L/6,
Length of open pipe = 5L/6
so, fundamental frequency = V/2(5L/6)
First overtone = f2 = V/(5L/6)
S0, f1/f2 = 5/8